### v6

# Universal framework for anisotropic particles with resonance laws and splitting

Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures. Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles. Closed-form solutions reveal axial permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory and highlight its predictive power. This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
Main
Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability to provide precise control over the spatial response In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [1, 2, 3, 4, 5, 6, 7]. While isotropic nanostructures have widespread use [8], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [9, 10]. When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [11], holding promise for a new generation of anisotropic resonators and biomarkers. In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [12, 13, 14, 15, 16, 17]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules [18, 19, 20, 21].
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime, Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [22, 23, 24, 25, 26, 27, 28]. Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Remarkably, to date, the only structure in physics for which closed-form quasistatic eigenmodes and eigenvalues have been derived is the subwavelength isotropic sphere [40, 41, 31]. Recently, the emergence of anisotropic materials in optics [9, 42, 43, 10, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53]. Moreover, uniaxial and biaxial materials have been shown to be very common in semiconductors and crystals, including the following: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and $\alpha$-molybdenum trioxide $(\alpha$-MoO3), with various geometries investigated [54, 10, 55, 16]. Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55]. However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior. These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems. This motivates our exploration of a new theoretical framework. The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [10, 56] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [57, 58]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [10, 59, 60, 61, 55].
Figure 1: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres. The modes of the isotropic sphere are directed along the field/dipole excitation direction [41, 31] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies). The high-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to superimposed modes.
[…]
Results
Our starting point in calculating the eigenstates of uniaxial and biaxial anisotropic spheres in the quasistatic regime is solving Laplace’s equation without a source, which reads for anisotropic media [67]
$\nabla\cdot\overleftrightarrow{\epsilon_{n}}\nabla\psi_{n}=0,$
where $\psi_{n}$ is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor $\overleftrightarrow{\epsilon_{n}}=\overleftrightarrow{\epsilon_{1n}}$ inside the inclusion and $\overleftrightarrow{\epsilon_{n}}=I\epsilon_{2}$ in the host medium ($I$ is the identity matrix). Note that the physical permittivity tensor $\overleftrightarrow{\epsilon_{1}}(\omega)$ generally depends on $\omega$ and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, $\boldsymbol{E}_{n}=\nabla\psi_{n}$ have to satisfy the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition. In addition to the challenge that the symmetry of the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely $\epsilon_{1i}\neq\epsilon_{1j}.$ As a result, the widely-employed spherical harmonics are in general not eigenstates of anisotropic spherical particles. In Fig. 1, left column, we present isotropic, uniaxial, and biaxial spheres, which are the most common types of spheres.
We now analyze the eigenmodes and eigenpermittivities of such anisotropic spheres. We start with the dipole response and proceed to the high-order mode responses. The anisotropic dipole eigenstates and eigenpermittivities of a sphere can be deduced from the scattering analyses in the literature [37] and read for a dipole mode oriented along $z$
$\tilde{\psi}_{l=1,m=0}=\frac{3}{4\pi}\frac{1}{\sqrt{a}}\left\{\begin{array}[]{cc}\frac{r}{a}\cos\theta&\,r<a\\ \frac{a^{2}}{r^{2}}\cos\theta&r\geq a\end{array}\right.,\,\,\,\epsilon_{1z}=-2,$ (3)
While these eigenstates and eigenpermittivities appear to be identical to the isotropic ones [37], here the modes are oriented along the resonant crystal axis independently of the incoming field polarization, even though the polarization direction does affect the excitation magnitude.
We now proceed to the high-order modes of anisotropic spheres. In our analysis we choose to express the eigenstates in cartesian coordinates due to the crystal symmetry with functions that are suitable for anisotropic medium i.e., ones that enable $\epsilon_{1i}\neq\epsilon_{1j}$ when substituted in Laplace’s equation and satisfy boundary conditions. Since the field outside the sphere can be expanded with spherical harmonics, the high-order spherical harmonics are valid for isotropic medium [31, 32], and there is continuity of the potential on the particle envelope, we suggest that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [41, 31]. To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at $\epsilon_{1ni}=\epsilon_{1nj}$ without having to satisfy $\epsilon_{1ni}=\epsilon_{1nk},$ so that the modes can be excited when $\epsilon_{1i}(\omega)=\epsilon_{1j}(\omega)\approx\epsilon_{1in}=\epsilon_{1jn}.$ While the isotropic sphere eigenmodes $\psi^{m}_{l}$ and their superpositions satisfy anisotropic Laplace’s equation for $\epsilon_{1x}=\epsilon_{1y}=\epsilon_{1z},$ when $\partial^{2}{\psi^{m}_{l}}/\partial x_{k}^{2}=0$ and $\partial^{2}{\psi^{m}_{l}}/\partial x_{i}^{2}\neq 0,\,\,\partial^{2}{\psi^{m}_{l}}/\partial x_{j}^{2}\neq 0,$ they naturally satisfy the above requirement. Interestingly, such functions can be composed of the isotropic-sphere eigenstates $\psi_{l}^{\pm{(l-1)}},\psi^{\pm{l}}_{l}.$ Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
$\tilde{\psi}_{l,u1,2}=\psi_{l}^{\pm\left(l-1\right)}=\left(x\pm iy\right)^{l-1}zf_{l}\left(r\right),$
$\left(l-1\right)\epsilon_{1x}+\epsilon_{1z}=-\epsilon_{2}\left(l+1\right),\,\,\,\epsilon_{1x}=\epsilon_{1y},$
$\tilde{\psi}_{l,u3,4}=\psi_{l}^{\pm l}=\left(x\pm iy\right)^{l}f_{l}\left(r\right),\,\,\epsilon_{1li}=\epsilon_{1lj}=-\epsilon_{2}\frac{l+1}{l},$
$f_{l}\left(r\right)=\left\{\begin{array}[]{cc}\frac{1}{\sqrt{la}}(1/a)^{l}&r<a\\ \frac{1}{\sqrt{la}}a^{l+1}/r^{2l+1}&r\geq a\end{array}\right$., (6)
where $x,y,z$ correspond to $x_{i},x_{j},x_{k}$ and the first condition holds if either $\epsilon_{x}=\epsilon_{y}<0$ or $\epsilon_{z}<0$, whereas the second holds only for $\epsilon_{x}=\epsilon_{y}<0,$ resulting in resonance splitting for $\epsilon_{x}=\epsilon_{y}<0$ (two closely-spaced resonances instead of one for isotropic spheres). One can see that since in the boundary conditions $\partial\psi^{\pm{l}}_{l}/\partial x_{k}=0$ and $\partial\psi^{\pm{(l-1})}_{l}/\partial x_{k}\neq 0,$ modes composed of $\psi^{\pm{l}}_{l}$ and $\psi_{l}^{\pm{(l-1)}}$ satisfy the isotropic eigenpermittivities [41, 31] for the equal permittivity axes and new eigenpermittivity relations, respectively. However, our modes in both cases are directional as opposed to the isotropic sphere modes. Alternatively, one could select the basis $\psi_{l}^{l}\pm\psi_{l}^{-l},\,\,\psi_{l}^{l-1}\pm\psi_{l}^{-(l-1)}$ for uniaxial spheres. Note also that the resonance conditions depend on $\epsilon_{2},$ which can be utilized to sense the environment in which the particle is situated.
[…]
In some cases, superpositions of the isotropic sphere modes satisfy $\partial^{2}\tilde{\psi_{n}}/\partial x_{k}^{2}=\partial^{2}\tilde{\psi_{n}}/\partial x_{i}^{2}=0,$ and anisotropic Laplace’s equation is valid for arbitrary values of $\epsilon_{1x},\,\epsilon_{1y},\,\epsilon_{1z}$, which have to obey only the boundary condition. Thus, we derive the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
$\tilde{\psi}_{2,b1}=x_{i}x_{j}f_{l}(r),\,\,\tilde{\psi}_{3,b}=x_{i}x_{j}x_{k}f_{l}(r),\,\tilde{\psi}_{4,b1}=x_{i}^{2}x_{j}x_{k}f_{l}(r),$
$\epsilon_{1i}+\epsilon_{1j}=-3,\,\sum_{p}\epsilon_{1p}=-4,\,\sum_{p}\epsilon_{1p}+\epsilon_{1i}=-5,\,\,i\neq j.$ (7)
These eigenstates are superpositions of spherical harmonics in all space, as is explained in Methods. Note that when imposing the boundary condition, this class of functions preserves the functional form under $\partial\psi_{bl}/\partial x_{i}\hat{r}^{i}$ inside the particle. Thus, the axial permittivities satisfy novel and interesting sum rules, this time for a general biaxial medium. These sum rules result in resonance splitting compared to the isotropic-sphere resonances: anisotropic bulk media usually exhibit a distinct resonance in each axis. Thus, for a resonance in a given axis in the bulk medium ($\epsilon_{1i}<0$), usually the sums result in two resonances occurring at different frequencies. Hence, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances. In the Supplementary Material (SM) we show that the derived uniaxial modes form a complete set and if additional biaxial modes exist they are typically negligible. In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. One can see that anisotropic particles support significantly different modes, which typically appear in several spectral regions as each axial permittivity experiences a resonanance at a different frequency (colors correspond to frequencies). In addition, their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere with the bulk permittivities of $\alpha-\mathrm{MoO_{3}}$. The scattering spectra and $|\mathbf{E}|^{2}$ of the eigenmodes on a $\alpha-\mathrm{MoO_{3}}$ sphere surface for the dipole (b), second-order (c), and third-order (d) modes. Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk. (e) $|\mathbf{E}|$ in the radiation patterns of two representative modes of $\psi=xz,\,\,\psi=xyz$ in the xyxy plane.
In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and $|\mathbf{E}|^{2}$ of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes, which also display a degeneracy. In Fig. 2 (e) we show $|\mathbf{E}|$ in the radiation patterns of the uniaxial sphere excited by an oscillating dipole (Methods). In Fig. 3 we present a biaxial sphere and the axial permittivities of $\alpha-\mathrm{MoO_{3}}$ bulk with three resonances. We then show the scattering spectra and $|\mathbf{E}|^{2}$ on the sphere surface for the dipole (b), second-order (c), and third-order modes (d). Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts. However, here the degeneracy is lifted since there are three different permittivities and the purely-uniaxial modes are not present. Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, in stark contrast to the situation in isotropic spheres. Fig. 3 (e) presents $|\mathbf{E}|$ in the radiation patterns of representative biaxial-sphere modes.
We now analyze the high-order modes of uniaxial and biaxial ellipsoids in analogy to the anisotropic spheres. Similarly, we suggest that the anisotropic ellipsoid modes are specific modes or superpositions of isotropic ellipsoid modes and consider the modes expressed in cartesian coordinates (Niven’s functions). We apply our conditions with the second derivatives to these modes and observe that while the first type of Niven’s functions do not satisfy them, the second type do satisfy them and inside the inclusion correspond to the biaxial sphere modes [68, 69, 70, 71]. To impose continuity of $D_{\perp}$ for $l=2$ we project $\boldsymbol{D}$ in the direction perpendicular to the interface with $v_{\perp}=\left(\frac{x}{a^{2}},\frac{y}{b^{2}},\frac{z}{c^{2}}\right)$ the unit vector and obtain for $\psi_{\mathrm{ins}}=xz$
$\mathbf{E}_{\mathrm{ins}}\propto z\hat{x}+x\hat{z},\,\,\tilde{D}_{\mathrm{ani}\,2\perp}^{2}=xz\left(\frac{\epsilon_{x}}{a^{2}}+\frac{\epsilon_{z}}{c^{2}}\right).$
Comparing $D_{\perp}$ inside the ellipsoid to the isotropic case $\tilde{D}_{\mathrm{iso}\,2r}^{2}=zx\epsilon_{1}\frac{1}{a^{2}}+xz\frac{1}{c^{2}}\epsilon_{1},$ we obtain that since $D_{\mathrm{ext}}$ and $E_{\parallel}$ (inside and outside) are the same for this isotropic and anisotropic cases, we can equate $D_{\perp\,\mathrm{in,iso}}=D_{\perp\,\mathrm{in,ani}}$ to obtain the general (sum rule) resonance condition:
$\frac{1}{a^{2}}\epsilon_{1x}+\frac{1}{c^{2}}\epsilon_{1z}=\left(\frac{1}{a^{2}}+\frac{1}{c^{2}}\right)\epsilon_{1xz\,\mathrm{iso}},$
where [68]
$\epsilon_{1xz\,\mathrm{iso}}=1-\left(\left(\frac{a_{x}a_{y}a_{z}}{2}\right)\left(a_{x}^{2}+a_{z}^{2}\right)I_{xz}\right)^{-1},$
$I_{\alpha\beta}=\intop_{0}^{\infty}\frac{du}{\left(u+a_{\alpha}^{2}\right)\left(u+a_{\beta}^{2}\right)R\left(u\right)}.$ (8)
For the sphere limit we get $\epsilon_{1x}+\epsilon_{1z}=-3$ as expected. The modes in the form $\tilde{\psi}\propto ax^{2}-by^{2},$ similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive (two conditions instead of one) and not likely to be realized experimentally. We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion $\tilde{\psi}_{3}=xyz.$ Similarly, we equate $D_{\perp}$ inside the ellipsoid in the isotropic and anisotropic cases and obtain the resonance condition:
$\left(\frac{\epsilon_{1x}}{a^{2}}+\frac{\epsilon_{1y}}{b^{2}}+\frac{\epsilon_{1z}}{c^{2}}\right)=\epsilon_{3\,\mathrm{iso}}\left(\frac{1}{a^{2}}+\frac{1}{b^{2}}+\frac{1}{c^{2}}\right),$
$\epsilon_{3\,\mathrm{iso}}=1-\left[\frac{\left(a_{1}a_{2}a_{3}\right)^{3}}{2}\left(\sum_{\alpha=1}^{3}a_{\alpha}^{-2}I_{123}\right)\right]^{-1},I_{123}=\intop_{0}^{\infty}\frac{du}{R^{3}\left(u\right)}.$
We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement $\epsilon_{1x}+\epsilon_{1y}+\epsilon_{1z}=3\epsilon_{3\,\mathrm{iso}}=-4.$ In this way one can readily find the fourth-order biaxial ellipsoid mode and resonance condition.
Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius $a=11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(1,0,1)\cdot 22\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$. Top: $\int|\mathbf{E}|^{2}da$ over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement. The circles are the peak widths obtained analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ compared to the ones calculated from the simulations, with very good agreement. (a),(b),(c),(d),(e) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=-2,$ $\epsilon_{1x}=-1.5,$ $\epsilon_{1x}+\epsilon_{1z}=-3,$ $\epsilon_{1x}=-4/3,$ $2\epsilon_{1x}+\epsilon_{1x}=-4$ calculated at f=46.301,f=46.301,\, $46.619,$  44.53,\,44.53,\,  46.725,\,46.725, $45.401\mathrm{(THz)},$ respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement. Bottom: $\int|\mathbf{E}|^{2}da$ over the particle envelope calculated in COMSOL for a dipole located at $\mathbf{r}_{0}=(1,0,0)31\mathrm{nm}$ (same dipole distance as the previous case) with a dipole moment of $\mathbf{P}=(1,0,0)\,\mathrm{A\cdot m}.$ As predicted the $m=\pm(l-1)$ modes disappeared from the spectrum.
Refer to caption Refer to caption
(a) (b)
Figure 5: Radiation pattern calculations for a uniaxial particle excited similarly to Fig. 4 (top). (a) Closed-form radiation-pattern plot of $|\mathbf{E}|$ corresponding to $\psi_{2}^{2}+\psi_{2}^{-2}$ outside the particle. (b) Full-wave COMSOL simulation of the scattered field for a uniaxial particle close to a resonance with $\epsilon_{1x}=-1.49$, $\epsilon_{1y}=-1.49$, $\epsilon_{1z}=1$ in $r>2\lambda$.
Figure 6: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 sphere of a radius $a=11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(18,18,18)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$. Top: $\int|\mathbf{E}|^{2}da$ calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement. (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=-2,$ $\epsilon_{1y}=-2,$ $\epsilon_{1x}+\epsilon_{1y}=-3,$ $\epsilon_{1x}+\epsilon_{1z}=-3,$ $\epsilon_{1y}+\epsilon_{1z}=-3,$ $\epsilon_{1x}+\epsilon_{1y}=-3,$ $\epsilon_{1x}+\epsilon_{1y}+\epsilon_{1z}=-4,$ calculated at $f=27.7,29.7,26.6,27,29.2,29.33,26.167\,\mathrm{(THz)}$ respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
Figure 7: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 ellipsoid with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(22,22,22)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$. Top: Surface integral over the particle envelope $\int|\mathbf{E}|^{2}da$ calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement. (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig. 5, defined in the ellipsoid volume).
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$. This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by $\left\langle\psi_{k}|\psi_{0}\right\rangle=\int\theta_{1}\nabla\psi_{k}^{*}\cdot\nabla\psi_{0}d\mathbf{r^{\prime}},$ where $\theta_{1}$ is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the $x$ axis. In this case, we do not expect the $m=\pm(l-1)$ modes to show in the spectrum since in order to excite them a $z$ component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$ compared to the surface integral of $|\mathbf{E}|^{2}$ over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to $\epsilon_{x}=-1.5,$ is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by $\omega/\Delta\omega.$ We then plot for all the observed peaks $|\mathbf{E}|$ inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot $\int|\mathbf{E}|^{2}da$ for a dipole located at the same distance as the previous case and oriented along $x$ ($\mathbf{r}_{0}=(1,0,0)$31nm). As we predicted the $m=\pm(l-1)$ modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode $\psi_{2}^{2}+\psi_{2}^{-2}$ with the resonance condition $\epsilon_{1x}=\epsilon_{1y}=-1.5$ outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with $R=46$ nm and $\epsilon_{1x}=-1.49,\epsilon_{1y}=-1.49,\epsilon_{1z}=1,$ excited by a current source similarly to Fig. 4 (top) in the radiation zone in the radiation zone $(r>2\lambda)$. Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
We then considered a biaxial subwavelength sphere composed of $\alpha$-MoO3 with the permittivity model in Ref. [54]. The sphere is excited by a dipole located at $\mathbf{r}_{0}=(18\,\mathrm{nm},18\,\mathrm{nm},18\,\mathrm{nm})$ with the dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$ We performed frequency domain simulations with a frequency sweep between $f=25.5\,\mathrm{(THz)}$ and $f=30.8\,\mathrm{(THz)}$ with 175 frequencies using COMSOL. Here, again, the dipole moment is directed along $\hat{r}$, to preserve the angular dependency of the mode in the mode-excitation coefficient. In Fig. 6 we present the result of the surface integral $\int|\mathbf{E}|^{2}da$ over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$. Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig. 6. We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes with excellent agreement. We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm},$ composed of $\alpha$-MoO3 with the permittivity model in Ref. [54]. The ellipsoid is excited by a dipole located at $\mathbf{r}_{0}=(22\,\mathrm{nm},22\,\mathrm{nm},22\,\mathrm{nm})$ with the dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$ We performed frequency domain simulations with a frequency sweep between $f=25.5\,\mathrm{(THz)}$ and $f=30.8\,\mathrm{(THz)}$ with 175 frequencies using COMSOL. In Fig. 7 we present the result of the surface integral $\int|\mathbf{E}|^{2}da$ over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$. We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes (equivalent to Fig. 6) with excellent agreement. We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: $\epsilon_{\mathrm{iso}\,x}=-1.387,$ $\epsilon_{\mathrm{iso}\,y}=-2.092,$ $\epsilon_{\mathrm{iso}\,xy}=-1.295,$ $\epsilon_{\mathrm{iso}\,yz}=-1.748,$ $\epsilon_{\mathrm{iso}\,xz}=-1.385,$ $\epsilon_{\mathrm{iso}\,xyz}=-1.287.$ As can be seen, for an isotropic ellipsoid with these semi axis ratios $\epsilon_{\mathrm{iso}\,x}\approx\epsilon_{\mathrm{iso}\,xz}$ and $\epsilon_{\mathrm{iso}\,xy}\approx\epsilon_{\mathrm{iso}\,xyz},$ and due to the anisotropic material these resonances are split, as can be seen in Fig. 7. In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental. Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
Discussion
We derived the modes and eigenpermittivity relations of uniaxial and biaxial particles in the quasistatic regime. We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting. We also identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes. We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry. Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [73, 74, 75, 76, 32]. The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59]. Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 77, 78]. Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, Y. Reches, and I. Kalyan are acknowledged for the insightful comments. We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
I Methods
I.1 Correspondence between the derived modes and the radiation patterns
Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres. We utilize the TM vector spherical harmonics (VSH) modes for a general spherical particle given by $\mathbf{E}\propto\nabla\times f_{l}\left(r\right)\boldsymbol{{X}}_{lm}\left(\theta,\phi\right)$ where
$f_{l}\left(r\right)=\left\{\begin{array}[]{cc}j_{l}\left(\sqrt{\epsilon_{1}}kr\right)&r<a\\ h_{l}^{\left(1\right)}\left(kr\right)&r>a\end{array}\right$.,
and $j_{l},h_{l}^{\left(1\right)}$ are the spherical bessel function and spherical Hankel function of the first kind and $\boldsymbol{X}_{lm}=\left(\boldsymbol{r}\times\nabla\right)Y_{lm}$ [79, 20]. Calculating the eigenpermittivity from these modes in the limit of $a\ll\lambda$ gives the quasistatic eigenvalues. For example, the full electrodynamic mode gives the following eigenpermittivities for a sphere with a=11.5a=11.5nm and $\lambda=600$nm : $\epsilon_{1,l=1}=-2.035-0.00357i,\epsilon_{1,l=2}=-1.505-2.12\cdot 10^{-6}i,\epsilon_{1,l=3}=-1.335-7.292\cdot 10^{-10}i,$ which are very close to the quasistatic results of $\epsilon_{1l}=-(l+1)/l.$ Even though these modes are usually considered for isotropic spheres, the results can be applied to anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs. We write a TM VSH outside the sphere:
$\nabla\times h_{l}^{\left(1\right)}\left(kr\right)\boldsymbol{X}_{lm}=\nabla h_{l}^{\left(1\right)}\times\boldsymbol{X}_{lm}+h_{l}^{\left(1\right)}\left(r\right)\nabla\times\boldsymbol{X}_{lm},$
Using the vector identity for $\boldsymbol{a}\times\boldsymbol{b}\times\boldsymbol{c}=\mathbf{b}(\mathbf{a}\cdot\mathbf{c})-\mathbf{c}(\mathbf{a}\cdot\mathbf{b})$ we get for the first term:
$\nabla h_{l}^{\left(1\right)}\times\boldsymbol{X}_{lm}=\nabla h_{l}^{\left(1\right)}\times\boldsymbol{r}\times\nabla Y_{lm}=-r\frac{\partial h_{l}^{\left(1\right)}(kr)}{\partial r}\nabla Y_{lm}.$
For the second term we arrive at
$h_{l}^{\left(1\right)}\left(kr\right)\nabla\times\boldsymbol{X}_{lm}=h_{l}^{\left(1\right)}\left(kr\right)\left(\boldsymbol{r}\nabla^{2}Y_{lm}-\nabla Y_{lm}\right),$
where we have used the vector identity for $\nabla\times\left(\boldsymbol{A}\times\boldsymbol{B}\right),$ $\nabla\cdot\boldsymbol{r}=3,$ $(\nabla Y_{lm}\cdot\nabla)\mathbf{r}=\nabla Y_{lm},$ $\left(\boldsymbol{r}\cdot\nabla\right)\nabla Y_{lm}=-\nabla Y_{lm}.$ Then, we utilize the spherical-harmonic identity [80]
$\nabla^{2}Y_{lm}=-l\left(l+1\right)/r^{2}Y_{lm},$
to obtain
$h_{l}^{\left(1\right)}\left(kr\right)\nabla\times\boldsymbol{X}_{lm}=h_{l}^{\left(1\right)}\left(kr\right)\left[-l\left(l+1\right)/rY_{lm}\hat{r}-\nabla Y_{lm}\right].$
All in all, we get:
$\nabla\times h_{l}^{\left(1\right)}\left(kr\right)\boldsymbol{X}_{lm}$
$=-r\frac{\partial h_{l}^{\left(1\right)}(kr)}{\partial r}\nabla Y_{lm}+h_{l}^{\left(1\right)}\left(kr\right)\left(-\frac{l\left(l+1\right)}{r}Y_{lm}\hat{r}-\nabla Y_{lm}\right).$
We see that when $r\rightarrow 0$ since $h_{l}^{\left(1\right)}\propto\frac{e^{ikr}}{r^{l+1}},$ $r\partial h_{l}^{\left(1\right)}/\partial r\propto\frac{e^{ikr}}{r^{l+1}}$ the overall scaling is of $1/r^{l+2},$ in agreement with the quasistatic analysis. For $r\rightarrow\infty$ since $\frac{\partial h_{l}^{(1)}}{\partial r}=\frac{ike^{ikr}}{r}-\frac{e^{ikr}}{r^{2}},\,\underset{r\to\infty}{\mathrm{lim}}\frac{\partial h_{l}^{(1)}}{\partial r}=\frac{ike^{ikr}}{r}$ and $\underset{r\to\infty}{\mathrm{lim}}\frac{h_{l}^{(1)}}{r}=\frac{e^{ikr}}{r^{2}}$ the first term, which is transverse and has $1/r$ scaling as expected, dominates. Thus, one can conclude from the superposition of $Y_{lm}$ that participate in the quasistatic anisotropic modes, that the radiation patterns are of the form:
$\mathbf{E}_{\mathrm{rad}}\approx\sum_{l^{\prime},m^{\prime}}-ike^{ikr}\nabla Y_{l^{\prime}m^{\prime}},$
where the sum is over the mode orders which participate in the anisotropic-sphere modes. Please note that $\nabla Y_{lm}\propto\frac{1}{r}$ which gives $1/r$ overall scaling of the radiated field and we get the following radiation pattern:
$\mathbf{E}_{\mathrm{rad}}\approx-\sum_{l^{\prime},m^{\prime}}\text{$\frac{ike^{ikr}}{r}\left(\mathbf{e}_{\theta}\frac{\partial}{\partial\theta},\mathbf{e}_{\phi}\frac{1}{\sin\theta}\frac{\partial}{\partial\phi}\right)Y_{l^{\prime}m^{\prime}}$}.$
One can readily see that the $l=1,m=0$ mode of a subwavelength sphere gives the dipole radiation pattern of $\mathbf{E}_{\mathrm{rad}}=\frac{e^{ikr}}{r}\sin\theta\mathbf{e}_{\theta}$. In addition, the radiation patterns of high-order anisotropic particle modes differ from the isotropic case. Note that while the uniaxial-sphere modes are identical to the l,l1l,l-1 isotropic sphere modes, in nearly all practical scenarios, their radiation patterns differ due to the fact that an excitation source will excite distinct superpositions of modes in the two cases.
I.2 Biaxial sphere modes as superpositions of spherical harmonics
We express the second-order biaxial modes as superpositions of spherical harmonics:
$\tilde{\psi}_{2,b1}=\left(\psi_{2}^{-1}+\psi_{2}^{1}\right)/2=xzf_{l}(r),$
$\tilde{\psi}_{2,b2}=\left(\psi_{2}^{1}-\psi_{2}^{-1}\right)/2i=yzf_{l}(r),$
$\tilde{\psi}_{2,b3}=\left(\psi_{2}^{2}-\psi_{2}^{-2}\right)/2i=xyf_{l}(r).$
We now proceed to the third and fourth-order biaxial modes. We continue along the lines of the second-order biaxial mode and write $\tilde{\psi}_{3}=\frac{\psi^{2}_{3}-\psi^{-2}_{3}}{2i}\propto xyz,$ $\tilde{\psi}_{b4}=\psi_{4}^{2}-\psi_{4}^{-2}+\frac{\psi_{4}^{4}-\psi_{4}^{-4}}{\sqrt{7}\cdot 4i}=xyz^{2}.$ Using the procedure above, it can be shown that $\epsilon_{1x}+\epsilon_{1y}+\epsilon_{1z}=-4$ and $\epsilon_{x}+\epsilon_{y}+2\epsilon_{z}=-5$ are the resonance conditions, respectively. Note that the modes $x^{2}yz,\,xy^{2}z$ and their resonance conditions readily follow.
I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
I.3.1 Theory
We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the main text.
Dipole mode
The resonance condition is $\epsilon_{x}+2=0$ and we therefore find the maximum of $1/|\epsilon_{x}+2|,$ which occurs for hBN [54] at $f=4.6265\cdot 10^{13}\left(\frac{1}{s}\right).$ The mode field inside the particle is given by $\mathbf{E}_{\mathrm{ins}}=\hat{x}$.
2nd order modes
For the first 2nd-order mode we get from the theory $\epsilon_{1x}+1.5=0,$ which for hBN approximately occurs at $f=4.6655\cdot 10^{13}\left(\frac{1}{s}\right).$ We add the degenerate modes, which are excited with equal strengths:
$\psi_{22}^{\pm}=\left(x\pm iy\right)^{2},\,\,\psi_{22}^{\pm}=x^{2}\pm ixy-y^{2},$
$\psi_{22}^{+}+\psi_{22}^{-}\propto x^{2}-y^{2},$
$\boldsymbol{E}_{22}^{+}+\boldsymbol{E}_{22}^{-}=2x\hat{x}-2y\hat{y},\,\,\left|\boldsymbol{E}_{22}^{+}+\boldsymbol{E}_{22}^{-}\right|\propto\sqrt{x^{2}+y^{2}}.$
second mode
The derived resonance condition reads
$\epsilon_{1x}+\epsilon_{1z}+3=0.$
For hBN this is approximately satisfied for $f=4.453\cdot 10^{13}\left(\mathrm{\frac{1}{s}}\right).$ Now we calculate the field by adding the degenerate modes:
$\psi_{21}^{\pm}=\left(x\pm iy\right)z,\,\,\psi_{21}^{+}+\psi_{21}^{-}\propto xz,$
$\boldsymbol{E}_{21}^{+}+\boldsymbol{E}_{21}^{-}=z\hat{x}+x\hat{z},\,\,\left|\boldsymbol{E}_{21}^{+}+\boldsymbol{E}_{21}^{-}\right|=\sqrt{z^{2}+x^{2}},$
$\left|\boldsymbol{E}_{21}^{+}\left(z=0\right)+\boldsymbol{E}_{21}^{-}\left(z=0\right)\right|=\left|x\right|.$
3rd order modes
First mode
The resonance condition reads:
$\epsilon_{1x}+4/3=0.$
For hBN this occurs at $f=4.6798\cdot 10^{13}\left(\frac{1}{s}\right)$ . Adding the degenerate modes we get
$\psi^{\pm}=\left(x\pm iy\right)^{3},$
$\psi^{\pm}=x^{3}\pm 2ix^{2}y-y^{2}x\pm iyx^{2}-2xy^{2}\mp iy^{3},$
$\mathbf{E}^{\pm}=\left(3x^{2}\pm 4ixy-y^{2}\pm 2iyx-2y^{2}\right)\hat{x}$
$+\left(\pm 2ix^{2}-2xy\pm ix^{2}-4xy\mp 3iy^{2}\right)\hat{y}$.
To simplify things, we look at the $yz$ plane
$\mathbf{E}^{\pm}\left(x=0\right)=-3y^{2}\hat{x}+\left(\pm 3iy^{2}\right)\hat{y}$,
$\mathbf{E}^{+}\left(x=0\right)+\mathbf{E}^{-}\left(x=0\right)\propto-3y^{2}\hat{x}$,
$\left|\mathbf{E}^{+}\left(x=0\right)+\mathbf{E}^{-}\left(x=0\right)\right|\propto 3y^{2}.$
Second mode
The resonance condition $2\epsilon_{1x}+\epsilon_{1x}+4=0$ is satisfied for hBN for $f=4.5409\,\mathrm{\left(\frac{1}{s}\right)}.$ Adding the modes we obtain:
$\psi_{31}^{\pm}=\left(x\pm iy\right)^{2}z,$
$\boldsymbol{E}_{31}^{\pm}=\left(x\pm iy\right)z\hat{x}\pm i\left(x\pm iy\right)z\hat{y}+\left(x\pm iy\right)z\hat{y}$,
$\psi_{31}^{+}+\psi_{31}^{-}\propto\left(x^{2}-y^{2}\right)z,$
$\boldsymbol{E}_{31}^{+}+\boldsymbol{E}_{31}^{-}=2xz\hat{x}-2yz\hat{y}+\left(x^{2}-y^{2}\right)\hat{z}$,
$\left|\boldsymbol{E}_{31}^{+}+\boldsymbol{E}_{31}^{-}\right|=\sqrt{\left(2xz\right)^{2}+\left(2yz\right)^{2}+\left(x^{2}-y^{2}\right)^{2}},$
$\left|\mathbf{E}\left(z=0\right)\right|=\left|x^{2}-y^{2}\right|.$
We plot these field magnitude distributions in Fig. 4.
One can express the eigenvalue in the denominator $s_{1xl}$ in terms of the physical parameter $\epsilon_{1z}$ using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs. 2 and 3). Note that the first sum includes only the $z$ dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper. It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.

### v7

@@
-Universal framework for anisotropic particles with resonance laws and splitting
+Universal framework for anisotropic nanoparticles with resonance laws and splitting
 Abstract
-Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures.
-Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
+Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and mode response in subwavelength structures.
+Anisotropic nanoparticles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
 Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles.
-Closed-form solutions reveal axial permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
+Closed-form solutions reveal axial-permittivity sum rules and material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
 Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss.
-Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory and highlight its predictive power.
-This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
+Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory.
+This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, enabling a new generation of photonic devices with tunable multispectral response and directional emission.
 Main
-Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability to provide precise control over the spatial response In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [1, 2, 3, 4, 5, 6, 7].
+Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, due to their ability to provide precise control over the response and produce strong, inherently passive behavior without the need for gain [1].
+In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [2, 3, 4, 5, 6, 7, 8].
 While isotropic nanostructures have widespread use [9], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [10, 11].
@@
 Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [30, 31, 32, 1, 33, 34, 35, 36, 37, 38, 39].
-Remarkably, to date, the only structure in physics for which closed-form quasistatic eigenmodes and eigenvalues have been derived is the subwavelength isotropic sphere [40, 41, 31].
+Remarkably, to date, the only structure in physics for which closed-form eigenmode response has been derived is the subwavelength isotropic sphere [40, 41, 32] (without residual one-dimensional integrals [30]).
 Recently, the emergence of anisotropic materials in optics [10, 42, 43, 11, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53].
-Moreover, uniaxial and biaxial materials have been shown to be very common in semiconductors and crystals, including the following: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and $\alpha$-molybdenum trioxide $(\alpha$-MoO3), with various geometries investigated [54, 10, 55, 16].
+In particular, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and $\alpha$-molybdenum trioxide $(\alpha$-MoO3), making them the most promising among anisotropic particles for a broad range of nanotechnology applications [54, 11, 55, 17].
 Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55].
-However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior.
+However, whereas previous approaches have successfully treated isotropic particles, fundamental challenges arise when analyzing nanostructures with cartesian anisotropy—particularly in predicting higher-order mode behavior—which hinders their theoretical understanding.
 These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
@@
 Results
-Our starting point in calculating the eigenstates of uniaxial and biaxial anisotropic spheres in the quasistatic regime is solving Laplace’s equation without a source, which reads for anisotropic media [67]
+To allow several degrees of freedom of the permittivities in the different axes, we derive the eigenstates [41, 32] of uniaxial and biaxial particles in the quasistatic regime.
+Our starting point is solving Laplace’s equation, which reads for anisotropic media [67]
 $\nabla\cdot\overleftrightarrow{\epsilon_{n}}\nabla\psi_{n}=0,$
@@
 Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition.
-In addition to the challenge that the symmetry of the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely $\epsilon_{1i}\neq\epsilon_{1j}.$ As a result, the widely-employed spherical harmonics are in general not eigenstates of anisotropic spherical particles.
 In Fig.
-1, left column, we present isotropic, uniaxial, and biaxial spheres, which are the most common types of spheres.
-We now analyze the eigenmodes and eigenpermittivities of such anisotropic spheres.
+1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most prevalent types of particles.
+We now derive the eigenmodes and eigenpermittivities of such anisotropic spheres.
 We start with the dipole response and proceed to the high-order mode responses.
-The anisotropic dipole eigenstates and eigenpermittivities of a sphere can be deduced from the scattering analyses in the literature [37] and read for a dipole mode oriented along $z$
+The anisotropic dipole eigenstates and eigenpermittivities can be deduced from the scattering analyses in the literature [37] and read for a dipole mode oriented along $z$
 $\tilde{\psi}_{l=1,m=0}=\frac{3}{4\pi}\frac{1}{\sqrt{a}}\left\{\begin{array}[]{cc}\frac{r}{a}\cos\theta&\,r<a\\ \frac{a^{2}}{r^{2}}\cos\theta&r\geq a\end{array}\right.,\,\,\,\epsilon_{1z}=-2,$ (3)
@@
 We now proceed to the high-order modes of anisotropic spheres.
-In our analysis we choose to express the eigenstates in cartesian coordinates due to the crystal symmetry with functions that are suitable for anisotropic medium i.e., ones that enable $\epsilon_{1i}\neq\epsilon_{1j}$ when substituted in Laplace’s equation and satisfy boundary conditions.
-Since the field outside the sphere can be expanded with spherical harmonics, the high-order spherical harmonics are valid for isotropic medium [31, 32], and there is continuity of the potential on the particle envelope, we suggest that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [41, 31].
-To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at $\epsilon_{1ni}=\epsilon_{1nj}$ without having to satisfy $\epsilon_{1ni}=\epsilon_{1nk},$ so that the modes can be excited when $\epsilon_{1i}(\omega)=\epsilon_{1j}(\omega)\approx\epsilon_{1in}=\epsilon_{1jn}.$ While the isotropic sphere eigenmodes $\psi^{m}_{l}$ and their superpositions satisfy anisotropic Laplace’s equation for $\epsilon_{1x}=\epsilon_{1y}=\epsilon_{1z},$ when $\partial^{2}{\psi^{m}_{l}}/\partial x_{k}^{2}=0$ and $\partial^{2}{\psi^{m}_{l}}/\partial x_{i}^{2}\neq 0,\,\,\partial^{2}{\psi^{m}_{l}}/\partial x_{j}^{2}\neq 0,$ they naturally satisfy the above requirement.
-Interestingly, such functions can be composed of the isotropic-sphere eigenstates $\psi_{l}^{\pm{(l-1)}},\psi^{\pm{l}}_{l}.$ Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
+The isotropic eigenstates that conform to the particle geometry (envelope boundary conditions and large-$r$ behavior) [41, 32] and the eigenstates satisfying cartesian anisotropic electrodynamic equation [48] are substantially different—making analytical solutions appear unattainable.
+Nevertheless, we identify two distinct classes of joint-property eigenfunctions.
+As we consider the geometric requirements to be the more stringent ones, we focus on the isotropic-particle eigenfunctions and impose additional conditions, to show that some of these eigenstates or their superpositions in fact exhibit both properties.
+Crucially, if these eigenstates satisfy $\partial^{2}{\psi^{m}_{l}}/\partial x_{k}^{2}=0$ inside the inclusion, they obey uniaxial Laplace’s equation inside the inclusion since $\nabla\cdot\overleftrightarrow{\epsilon_{n}}\nabla\psi^{m}_{l}=\epsilon_{1n}\nabla^{2}\psi^{m}_{l}=0.$ Strikingly, such functions are the isotropic-sphere eigenstates $\psi_{l}^{\pm{(l-1)}},\psi^{\pm{l}}_{l}$ and we obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
 $\tilde{\psi}_{l,u1,2}=\psi_{l}^{\pm\left(l-1\right)}=\left(x\pm iy\right)^{l-1}zf_{l}\left(r\right),$
 $\left(l-1\right)\epsilon_{1x}+\epsilon_{1z}=-\epsilon_{2}\left(l+1\right),\,\,\,\epsilon_{1x}=\epsilon_{1y},$
-$\tilde{\psi}_{l,u3,4}=\psi_{l}^{\pm l}=\left(x\pm iy\right)^{l}f_{l}\left(r\right),\,\,\epsilon_{1li}=\epsilon_{1lj}=-\epsilon_{2}\frac{l+1}{l},$
+$\tilde{\psi}_{l,u3,4}=\psi_{l}^{\pm l}=\left(x\pm iy\right)^{l}f_{l}\left(r\right),\,\,\epsilon_{1x}=\epsilon_{1y}=-\epsilon_{2}\frac{l+1}{l},$
 $f_{l}\left(r\right)=\left\{\begin{array}[]{cc}\frac{1}{\sqrt{la}}(1/a)^{l}&r<a\\ \frac{1}{\sqrt{la}}a^{l+1}/r^{2l+1}&r\geq a\end{array}\right$., (6)
-where $x,y,z$ correspond to $x_{i},x_{j},x_{k}$ and the first condition holds if either $\epsilon_{x}=\epsilon_{y}<0$ or $\epsilon_{z}<0$, whereas the second holds only for $\epsilon_{x}=\epsilon_{y}<0,$ resulting in resonance splitting for $\epsilon_{x}=\epsilon_{y}<0$ (two closely-spaced resonances instead of one for isotropic spheres).
-One can see that since in the boundary conditions $\partial\psi^{\pm{l}}_{l}/\partial x_{k}=0$ and $\partial\psi^{\pm{(l-1})}_{l}/\partial x_{k}\neq 0,$ modes composed of $\psi^{\pm{l}}_{l}$ and $\psi_{l}^{\pm{(l-1)}}$ satisfy the isotropic eigenpermittivities [41, 31] for the equal permittivity axes and new eigenpermittivity relations, respectively.
-However, our modes in both cases are directional as opposed to the isotropic sphere modes.
-Alternatively, one could select the basis $\psi_{l}^{l}\pm\psi_{l}^{-l},\,\,\psi_{l}^{l-1}\pm\psi_{l}^{-(l-1)}$ for uniaxial spheres.
+where $x,y,z$ correspond to $x_{i},x_{j},x_{k}.$ Uniaxial bulk exhibits two resonances with $\epsilon_{1x}=\epsilon_{1y}<0$ and $\epsilon_{1z}<0,$ and one can see that the first condition holds if either $\epsilon_{1x}=\epsilon_{1y}<0$ or $\epsilon_{1z}<0$, whereas the second holds only for $\epsilon_{1x}=\epsilon_{1y}<0.$ This results in resonance splitting for $\epsilon_{1x}=\epsilon_{1y}<0$ since there are two closely-spaced resonances instead of one for isotropic spheres.
+Our modes in both cases are directional as opposed to the isotropic sphere modes.
 Note also that the resonance conditions depend on $\epsilon_{2},$ which can be utilized to sense the environment in which the particle is situated.
 […]
-In some cases, superpositions of the isotropic sphere modes satisfy $\partial^{2}\tilde{\psi_{n}}/\partial x_{k}^{2}=\partial^{2}\tilde{\psi_{n}}/\partial x_{i}^{2}=0,$ and anisotropic Laplace’s equation is valid for arbitrary values of $\epsilon_{1x},\,\epsilon_{1y},\,\epsilon_{1z}$, which have to obey only the boundary condition.
+Importantly, some superpositions of the isotropic sphere modes satisfy $\partial^{2}\tilde{\psi_{n}}/\partial x_{k}^{2}=\partial^{2}\tilde{\psi_{n}}/\partial x_{i}^{2}=0$ inside the particle, and biaxial Laplace’s equation is valid for arbitrary values of $\epsilon_{1x},\,\epsilon_{1y},\,\epsilon_{1z}$, and they have to obey only the boundary condition.
 Thus, we derive the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
 $\tilde{\psi}_{2,b1}=x_{i}x_{j}f_{l}(r),\,\,\tilde{\psi}_{3,b}=x_{i}x_{j}x_{k}f_{l}(r),\,\tilde{\psi}_{4,b1}=x_{i}^{2}x_{j}x_{k}f_{l}(r),$
-$\epsilon_{1i}+\epsilon_{1j}=-3,\,\sum_{p}\epsilon_{1p}=-4,\,\sum_{p}\epsilon_{1p}+\epsilon_{1i}=-5,\,\,i\neq j.$ (7)
+$\epsilon_{1i}+\epsilon_{1j}=-3,\,\sum_{p}\epsilon_{1p}=-4,\,\sum_{p}\epsilon_{1p}+\epsilon_{1i}=-5.$ (7)
 These eigenstates are superpositions of spherical harmonics in all space, as is explained in Methods.
 Note that when imposing the boundary condition, this class of functions preserves the functional form under $\partial\psi_{bl}/\partial x_{i}\hat{r}^{i}$ inside the particle.
-Thus, the axial permittivities satisfy novel and interesting sum rules, this time for a general biaxial medium.
-These sum rules result in resonance splitting compared to the isotropic-sphere resonances: anisotropic bulk media usually exhibit a distinct resonance in each axis.
-Thus, for a resonance in a given axis in the bulk medium ($\epsilon_{1i}<0$), usually the sums result in two resonances occurring at different frequencies.
-Hence, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances.
-In the Supplementary Material (SM) we show that the derived uniaxial modes form a complete set and if additional biaxial modes exist they are typically negligible.
+These sum rules also result in resonance splitting compared to the isotropic-sphere resonances: for a resonance in a given axis in the bulk medium ($\epsilon_{1i}<0$), the sum rules typically result in two resonances occurring at distinct frequencies.
+In the Supplementary Material (SM) we derive eigenstate field expansions for any excitation source and show that the derived uniaxial modes form a complete set and if additional biaxial modes exist they are negligible.
 In Fig.
 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
-One can see that anisotropic particles support significantly different modes, which typically appear in several spectral regions as each axial permittivity experiences a resonanance at a different frequency (colors correspond to frequencies).
+One can see that anisotropic particles support significantly different modes, which appear in several spectral regions as each axial permittivity experiences a resonance at a different frequency (colors correspond to frequencies).
 In addition, their direction is determined only by the crystal orientation.
 Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
+We generalize this eigenmode derivation to the full-wave regime and obtain the Q factors analytically in Methods.
 Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle.
@@
 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances.
-We then show for a uniaxial sphere composed of hBN the spectrum and $|\mathbf{E}|^{2}$ of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes.
-Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes, which also display a degeneracy.
-In Fig.
-2 (e) we show $|\mathbf{E}|$ in the radiation patterns of the uniaxial sphere excited by an oscillating dipole (Methods).
+We then show for a uniaxial sphere composed of hBN the spectrum and $|\mathbf{E}|^{2}$ on the external surface of the sphere similarly to near-field measurements from the derived resonance conditions and modes, for the dipole (b), second-order (c), and third-order (d) modes.
+Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes (green rectangles), which also display a degeneracy.
+Fig.
+2 (e) shows $|\mathbf{E}|$ in the radiation patterns for an oscillating dipole excitation (Methods), which are significantly different from those of an isotropic sphere [21].
 In Fig.
 3 we present a biaxial sphere and the axial permittivities of $\alpha-\mathrm{MoO_{3}}$ bulk with three resonances.
-We then show the scattering spectra and $|\mathbf{E}|^{2}$ on the sphere surface for the dipole (b), second-order (c), and third-order modes (d).
+We then show the calculated scattering spectra and $|\mathbf{E}|^{2}$ on the sphere surface for the dipole (b), second-order (c), and third-order modes (d).
 Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts.
@@
 3 (e) presents $|\mathbf{E}|$ in the radiation patterns of representative biaxial-sphere modes.
-We now analyze the high-order modes of uniaxial and biaxial ellipsoids in analogy to the anisotropic spheres.
-Similarly, we suggest that the anisotropic ellipsoid modes are specific modes or superpositions of isotropic ellipsoid modes and consider the modes expressed in cartesian coordinates (Niven’s functions).
-We apply our conditions with the second derivatives to these modes and observe that while the first type of Niven’s functions do not satisfy them, the second type do satisfy them and inside the inclusion correspond to the biaxial sphere modes [68, 69, 70, 71].
-To impose continuity of $D_{\perp}$ for $l=2$ we project $\boldsymbol{D}$ in the direction perpendicular to the interface with $v_{\perp}=\left(\frac{x}{a^{2}},\frac{y}{b^{2}},\frac{z}{c^{2}}\right)$ the unit vector and obtain for $\psi_{\mathrm{ins}}=xz$
-$\mathbf{E}_{\mathrm{ins}}\propto z\hat{x}+x\hat{z},\,\,\tilde{D}_{\mathrm{ani}\,2\perp}^{2}=xz\left(\frac{\epsilon_{x}}{a^{2}}+\frac{\epsilon_{z}}{c^{2}}\right).$
-Comparing $D_{\perp}$ inside the ellipsoid to the isotropic case $\tilde{D}_{\mathrm{iso}\,2r}^{2}=zx\epsilon_{1}\frac{1}{a^{2}}+xz\frac{1}{c^{2}}\epsilon_{1},$ we obtain that since $D_{\mathrm{ext}}$ and $E_{\parallel}$ (inside and outside) are the same for this isotropic and anisotropic cases, we can equate $D_{\perp\,\mathrm{in,iso}}=D_{\perp\,\mathrm{in,ani}}$ to obtain the general (sum rule) resonance condition:
-$\frac{1}{a^{2}}\epsilon_{1x}+\frac{1}{c^{2}}\epsilon_{1z}=\left(\frac{1}{a^{2}}+\frac{1}{c^{2}}\right)\epsilon_{1xz\,\mathrm{iso}},$
-where [68]
-$\epsilon_{1xz\,\mathrm{iso}}=1-\left(\left(\frac{a_{x}a_{y}a_{z}}{2}\right)\left(a_{x}^{2}+a_{z}^{2}\right)I_{xz}\right)^{-1},$
+We now generalize our results to uniaxial and biaxial ellipsoids to enable geometric tunability of the response.
+Similarly, we suggest that the anisotropic ellipsoid modes are specific modes of isotropic ellipsoid modes in Niven function basis [68, 69, 70, 71].
+Applying our conditions involving the second derivatives of the potential, we find that the first type of Niven functions fails to satisfy them, whereas the second type does, and corresponds to the modes of the biaxial sphere inside the inclusion.
+We impose continuity of $D_{\perp}$ for the $l=2$ modes with $x_{i}x_{j}$ inside the inclusion by projecting $\boldsymbol{D}$ in the direction perpendicular to the interface and obtain generalized sum-rule resonance condition:
+$a_{i}^{-2}\epsilon_{1i}+a_{j}^{-2}\epsilon_{1j}=\left(a_{i}^{-2}+a_{j}^{-2}\right)\epsilon_{1ij}^{\mathrm{iso}}.$
+where $a_{i}$ are the ellipsoid semiaxes, and [68]
+$\epsilon^{\mathrm{iso}}_{1ij}=1-\left[\left(\frac{a_{i}a_{j}a_{k}}{2}\right)\left(a_{i}^{2}+a_{j}^{2}\right)I_{ij}\right]^{-1},$
 $I_{\alpha\beta}=\intop_{0}^{\infty}\frac{du}{\left(u+a_{\alpha}^{2}\right)\left(u+a_{\beta}^{2}\right)R\left(u\right)}.$ (8)
-For the sphere limit we get $\epsilon_{1x}+\epsilon_{1z}=-3$ as expected.
-The modes in the form $\tilde{\psi}\propto ax^{2}-by^{2},$ similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive (two conditions instead of one) and not likely to be realized experimentally.
+For the sphere limit we get $\epsilon_{1i}+\epsilon_{1j}=-3$ as expected.
 We proceed to the third-order mode of an anisotropic ellipsoid.
 We write the potential and field inside the inclusion $\tilde{\psi}_{3}=xyz.$ Similarly, we equate $D_{\perp}$ inside the ellipsoid in the isotropic and anisotropic cases and obtain the resonance condition:
-$\left(\frac{\epsilon_{1x}}{a^{2}}+\frac{\epsilon_{1y}}{b^{2}}+\frac{\epsilon_{1z}}{c^{2}}\right)=\epsilon_{3\,\mathrm{iso}}\left(\frac{1}{a^{2}}+\frac{1}{b^{2}}+\frac{1}{c^{2}}\right),$
-$\epsilon_{3\,\mathrm{iso}}=1-\left[\frac{\left(a_{1}a_{2}a_{3}\right)^{3}}{2}\left(\sum_{\alpha=1}^{3}a_{\alpha}^{-2}I_{123}\right)\right]^{-1},I_{123}=\intop_{0}^{\infty}\frac{du}{R^{3}\left(u\right)}.$
+$a_{i}^{-2}\epsilon_{1i}+a_{j}^{-2}\epsilon_{1j}+a_{k}^{-2}\epsilon_{1k}=(a_{i}^{-2}+a_{j}^{-2}+a_{k}^{-2})\,\epsilon_{3}^{\mathrm{iso}}.$ (9)
+$\epsilon_{3}^{\mathrm{iso}}=1-\left[\frac{\left(a_{1}a_{2}a_{3}\right)^{3}}{2}\left(\sum_{\alpha=1}^{3}a_{\alpha}^{-2}I_{123}\right)\right]^{-1},I_{123}=\intop_{0}^{\infty}\frac{du}{R^{3}\left(u\right)}.$
 We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement $\epsilon_{1x}+\epsilon_{1y}+\epsilon_{1z}=3\epsilon_{3\,\mathrm{iso}}=-4.$ In this way one can readily find the fourth-order biaxial ellipsoid mode and resonance condition.
+Finally, we verified our theory for uniaxial and biaxial spheres and a biaxial ellipsoid in extensive full-wave simulations with excellent agreement for the modes and the spectra (Methods).
+In our recent work in Ref.
+[55] we also verified the second-order mode response experimentally for two biaxial ellipsoids with hyperspectral mid-IR near-field imaging.
+Conclusion
+We derived the modes and eigenpermittivity relations of uniaxial and biaxial nanoparticles in the quasistatic regime as well as their radiation patterns.
+We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting.
+We also identified eigenfrequency degeneracy for uniaxial spheres, resulting in superimposed modes.
+Interestingly, as opposed to isotropic-material particles, high-order resonance frequencies in anisotropic particles are often lower than the dipole resonance.
+We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry.
+Finally, we derived the Q factors analytically and performed extensive full-wave simulations that highlighted the predictive power of the theory.
+Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [72, 73, 74, 75, 1].
+The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers [59] with enhanced specificity of the spectral signature.
+In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59].
+Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 76, 77].
+Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules.
+Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
+Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
+Methods
+.1 Full-wave simulations of uniaxial and biaxial particles
 Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius $a=11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(1,0,1)\cdot 22\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$.
@@
 (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=-2,$ $\epsilon_{1y}=-2,$ $\epsilon_{1x}+\epsilon_{1y}=-3,$ $\epsilon_{1x}+\epsilon_{1z}=-3,$ $\epsilon_{1y}+\epsilon_{1z}=-3,$ $\epsilon_{1x}+\epsilon_{1y}=-3,$ $\epsilon_{1x}+\epsilon_{1y}+\epsilon_{1z}=-4,$ calculated at $f=27.7,29.7,26.6,27,29.2,29.33,26.167\,\mathrm{(THz)}$ respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
-Figure 7: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 ellipsoid with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(22,22,22)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$.
+Figure 7: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 ellipsoid with semiaxes $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(22,22,22)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$.
 Top: Surface integral over the particle envelope $\int|\mathbf{E}|^{2}da$ calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement.
-(a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig.
-5, defined in the ellipsoid volume).
+(a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig.
+6, defined in the ellipsoid volume).
 Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations.
@@
 The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [67, 78, 32], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$.
-This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section.
+This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in Methods D.
 Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes.
@@
 We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
-Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm},$ composed of $\alpha$-MoO3 with the permittivity model in Ref.
+Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with semiaxes $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm},$ composed of $\alpha$-MoO3 with the permittivity model in Ref.
 [54].
@@
 Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
-Discussion
-We derived the modes and eigenpermittivity relations of uniaxial and biaxial particles in the quasistatic regime.
-We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting.
-We also identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes.
-We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry.
-Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [73, 74, 75, 76, 32].
-The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature.
-In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59].
-Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 77, 78].
-Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules.
-Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
-Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
-Acknowledgement
-E.
-E.
-Narimanov, A.
-Niv, Y.
-Mazor, J.
-Jeffet, Y.
-Reches, and I.
-Kalyan are acknowledged for the insightful comments.
-We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
-I Methods
-I.1 Correspondence between the derived modes and the radiation patterns
+.2 Correspondence between the derived modes and the radiation patterns
 Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres.
@@
 Note that while the uniaxial-sphere modes are identical to the l,l1l,l-1 isotropic sphere modes, in nearly all practical scenarios, their radiation patterns differ due to the fact that an excitation source will excite distinct superpositions of modes in the two cases.
-I.2 Biaxial sphere modes as superpositions of spherical harmonics
+.3 Biaxial sphere modes as superpositions of spherical harmonics
 We express the second-order biaxial modes as superpositions of spherical harmonics:
@@
 Note that the modes $x^{2}yz,\,xy^{2}z$ and their resonance conditions readily follow.
-I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
-I.3.1 Theory
+.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
+.4.1 Theory
 We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the main text.
@@
 It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
+Acknowledgement
+E.
+E.
+Narimanov, A.
+Niv, Y.
+Mazor, J.
+Jeffet, Y.
+Reches, and I.
+Kalyan are acknowledged for the insightful comments.
+We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.

### v8

@@
-Universal framework for anisotropic nanoparticles with resonance laws and splitting
+Universal framework for uniaxial and biaxial particles with resonance laws and splitting
 Abstract
 Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and mode response in subwavelength structures.
-Anisotropic nanoparticles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
-Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles.
+Uniaxial and biaxial particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
+Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of such nanoparticles.
 Closed-form solutions reveal axial-permittivity sum rules and material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
 Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss.
-Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory.
-This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, enabling a new generation of photonic devices with tunable multispectral response and directional emission.
-Main
+Full-wave simulations of h-BN and $\alpha$-MoO3 nanoparticles are consistent with the theoretical predictions and with our recently reported experimental observations.
+This framework unifies the understanding of anisotropic nanostructures across optics, magnetism, and thermal transport enabling a new generation of photonic devices with tunable multispectral response and directional emission.
+Introduction
 Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, due to their ability to provide precise control over the response and produce strong, inherently passive behavior without the need for gain [1].
@@
 Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55].
-However, whereas previous approaches have successfully treated isotropic particles, fundamental challenges arise when analyzing nanostructures with cartesian anisotropy—particularly in predicting higher-order mode behavior—which hinders their theoretical understanding.
-These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
+However, whereas previous approaches have successfully treated isotropic particles, fundamental challenges arise when analyzing nanostructures with cartesian anisotropy—particularly in predicting higher-order mode behavior—hindering their theoretical understanding.
+These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, the distinct differential equations in each medium, and the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
 This motivates our exploration of a new theoretical framework.
 The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [11, 56] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [57, 58].
-Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [11, 59, 60, 61, 55].
+Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [11, 59, 60, 55].
+Figure 1: Scope of applicability of the proposed framework
 Figure 2: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres.
@@
 Results
-To allow several degrees of freedom of the permittivities in the different axes, we derive the eigenstates [41, 32] of uniaxial and biaxial particles in the quasistatic regime.
+Closed-form eigenmodes and resonance conditions in uniaxial and biaxial spheres
+To allow several degrees of freedom of the permittivities in the different axes, we utilize the definition of an eigenstate [41, 32] as a source-free electric potential in the quasistatic regime applied here to uniaxial and biaxial particles.
 Our starting point is solving Laplace’s equation, which reads for anisotropic media [66]
@@
 Clearly, $\boldsymbol{E}_{n}=\nabla\psi_{n}$ have to satisfy the boundary conditions at the particle interface.
-Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition.
+Once anisotropic Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition.
 In Fig.
-1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most prevalent types of particles.
+2, left column, we present isotropic, uniaxial, and biaxial particles, which are the most prevalent types of particles.
 We now derive the eigenmodes and eigenpermittivities of such anisotropic spheres.
@@
 4 (e) presents $|\mathbf{E}|$ in the radiation patterns of representative biaxial-sphere modes.
+Analytic eigenmodes and resonance conditions in uniaxial and biaxial ellipsoids
 We now generalize our results to uniaxial and biaxial ellipsoids to enable geometric tunability of the response.
@@
 [55] we also verified the second-order mode response experimentally for two biaxial ellipsoids with hyperspectral mid-IR near-field imaging.
-Conclusion
+Discussion
 We derived the modes and eigenpermittivity relations of uniaxial and biaxial nanoparticles in the quasistatic regime as well as their radiation patterns.
@@
 Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [71, 72, 73, 74, 1].
-The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers [59] with enhanced specificity of the spectral signature.
-In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59].
+The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature.
+In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection.
 Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [59, 75, 76].
@@
 Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
-Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
+Finally, our results apply also to other fields of physics such as magnetism and heat conduction [56].
 Methods
@@
 (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig.
-6, defined in the ellipsoid volume).
-Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations.
+7, defined in the ellipsoid volume).
+We demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations.
 While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors.
 We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs.
-2,3 [54].
+3,4 [54].
 The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations.
@@
 Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig.
-6.
+7.
 We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes with excellent agreement.
@@
 We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes (equivalent to Fig.
-6) with excellent agreement.
+7) with excellent agreement.
 We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: $\epsilon_{\mathrm{iso}\,x}=-1.387,$ $\epsilon_{\mathrm{iso}\,y}=-2.092,$ $\epsilon_{\mathrm{iso}\,xy}=-1.295,$ $\epsilon_{\mathrm{iso}\,yz}=-1.748,$ $\epsilon_{\mathrm{iso}\,xz}=-1.385,$ $\epsilon_{\mathrm{iso}\,xyz}=-1.287.$ As can be seen, for an isotropic ellipsoid with these semi axis ratios $\epsilon_{\mathrm{iso}\,x}\approx\epsilon_{\mathrm{iso}\,xz}$ and $\epsilon_{\mathrm{iso}\,xy}\approx\epsilon_{\mathrm{iso}\,xyz},$ and due to the anisotropic material these resonances are split, as can be seen in Fig.
-7.
+8.
 In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental.
@@
 One can express the eigenvalue in the denominator $s_{1xl}$ in terms of the physical parameter $\epsilon_{1z}$ using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs.
-2 and 3).
+3 and 4).
 Note that the first sum includes only the $z$ dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper.